# Multi-scale phylodynamic modelling of rapid punctuated pathogen evolution

### v2

@@ A.5 Intra-host (within-host) selective pressure
A virus that has less opportunity to develop within-host diversity will be less effective at exploring the fitness landscape. We emulate this by biasing the fitness of mutant sequences generated from a parent sequence. At each time step where we need to generate mutations in the viral sequence associated with an infected host, we first independently generate $M$ candidate sequences and rank them by their fitness. We then sample one sequence from the top $X\leq M$ candidates to be the new sequence. The difference $M-X+1$ specifies the selective pressure strength. For example, if $M=X$, this is equivalent to generating only a single candidate sequence, indicating a weak selective pressure. On the other hand, if $M=100$ and $X=1$, then the selected sequence will be the fittest of 100 randomly generated candidate sequences, indicating a strong selective pressure and bias towards higher fitness. In other words, small values of $X$ imply more efficient exploration of the fitness landscape as might be expected during a chronic infection. This mechanism emulates within-host dynamics that are not explicitly modelled, since each infected agent is associated with only a single viral genome at each simulation cycle.
@@
    Randomly generate $M-X+1$ mutated genome sequences by independently selecting and changing $L$ positions in sequence $s$ for each, by substituting the corresponding nucleotide number by a uniformly sampled number between 00 and $3$. These generated mutated genome sequences become candidates for the upcoming selection process.
@@ Appendix B Artificial agent-based population
Prior to the simulation, we stochastically generated three distinct artificial populations of anonymous agents (corresponding to layer (B) Demographic shown in Fig. 12), used to examine evolutionary and epidemiological dynamics:
@@
We selected these population sizes for two reasons: (i) to investigate the relationship between the population size and phylodynamics, and (ii) to ensure computational feasibility. These considerations are based on the observation that the SARS-CoV-2 variants of concern may have emerged in countries of different population sizes (e.g., South Africa, with a population of approximately 59 million; UK, with a population of around 56 million; and Botswana, with a population over 2.4 million). In addition, our preliminary test simulations revealed that the computational cost increases non-linearly with the population size (see Section H). In this study, we focus on revealing the representative evolutionary characteristics of SARS-CoV-2, while maintaining computational efficiency by using the three populations listed above. However, given sufficient computational resources, much larger populations (e.g., over one billion population representing countries such as India) may also be generated in order to quantify the impact of population size on phylodynamic and disease dynamics.
The constructed artificial population captures essential demographic characteristics and commuting patterns represented in the latest Australian census and other datasets, as well as agent-to-agent interactions in various social mixing contexts, thus representing the population heterogeneity. Specifically, we generated the artificial population using demographic and travel data sourced from the Australian Bureau of Statistics (ABS) 2021 Census [32], international air traffic reports (detailing incoming passenger flows at Australian airports) [55], and educational registration records (including data on schools and students) [56]. Each agent in the artificial population was assigned multiple demographic attributes, such as age, gender, and residency location, alongside social mixing contexts across various settings: residential (e.g., household, household cluster, neighbourhood and statistical area which maps to a local government area), educational (e.g., classroom/school for agents aged 18 years or younger), and workplace (for agents aged over 18 years). Figure 5 provides a visual representation of the considered social mixing contexts. For each agent, the residential contexts are determined based on residential demographics, while the workplace and educational contexts are assigned based on commuting patterns (commute to work and class/school, depending on the agent’s age group).
@@ C.1 Multi-strain transmission model
The multi-strain transmission model of PhASE TraCE stochastically simulates pathogen transmission between hosts (Epidemic Layer shown in Fig. 12). This model introduced several new features to our previous ABM, AMTraC-19 [6, 23, 26, 24]. In this section, we describe how transmission is affected by the genome model described above.
@@ C.1.1 Susceptible-infectious transition
where $K(s_{j})$ represents the transmissibility of pathogen variant $s_{j}$, proportional to the corresponding basic reproductive number $R_{0}$, i.e., $K(s_{j})$ is the fitness of genome $s_{j}$ carried by agent $j$, as defined by equation (7); $n_{j}$ is the time cycle when agent $j$ started experiencing the onset of infection; and the function $f_{j}(n-n_{j})$ determines agent $j$’s infectivity over time according to the natural history model. At the time cycle $n$, agent $j$’s infectivity is $0<f_{j}(n-n_{j})\leq 1$ with the infectivity peak equal to 1. For uninfected agents, $f(\cdot)=0$. For simplicity, we assume that all variants share the same progression of disease, but differ in their transmissibility (determined by fitness $K$), and that only one variant can be transmitted during the infection process.
@@
where $G_{i}(n)$ is the set of social mixing contexts where interactions between agent $i$ and other agents may occur (depending on factors such as weekday or weekend, daytime or nighttime), and $A_{g}\backslash\{i\}$ is the list of all agents in the context $g\in G_{i}(n)$ except agent $i$. We shall expand Equation 9 in later sections to incorporate the effects of non-pharmaceutical interventions (Equation 11) and effects of immunity (Equation 20).
Eq. 9 defines the infection probability for agent $i$ interacting with other infected agents across all of its social contexts. However, this probability is an aggregate of the interactions and does not identify the specific source of infection, which must be known in order to copy its viral sequence to the newly infected host. We assign the source of infection by constructing a discrete distribution to randomly sample one agent from all potential infected agents $j$ across all social contexts $g\in G_{i}(n)$ to which agent $i$ belongs. The probability of selecting agent $j$ is given by
@@
where $\sigma_{i}=\sigma^{a}$ for adults ($\text{age}>18$) and $\sigma_{i}=\sigma^{c}$ for children ($\text{age}\leq 18$). The asymptomatic infectious agents have lower infectivity (i.e., lower $f_{i}(n-n_{i})$, compared to the symptomatic agents, which also affects the strength of transmission to other susceptible agents.
@@ C.1.2 Infectious-recovered and recovered-susceptible transitions
Following the natural history model, an infectious agent recovers after a certain recovery period dependent on the host category (see Section G). We assume that both typical and chronically infected hosts develop infectivity fairly quickly from the onset of infection (following a lognormal distribution). However, chronically infected hosts experience a significantly longer recovery than the typical infected hosts. The natural history model describing the disease progression for both host categories is illustrated in Fig. 16.
@@ C.2 Non-Pharmaceutical Interventions
The assignment of NPI-compliant agents is determined based on a Bernoulli process with the probability specified by the macro-distancing parameters. While an agent may comply with multiple NPIs, their interaction strengths, $F_{g}(j)$, can only be adjusted to one NPI. We therefore use the parameters for only the first NPI to which an agent complies in the following ordered list: CI, HQ, SD, and SC. The micro and macro-distancing parameters of CI vary depending on the host type, where chronically infected hosts are fully compliant (i.e., macro-distancing level is set to $1.0$), with significantly reduced interaction strengths across all social contexts (Tables 1 and 2). This parametrisation agrees with the public health recommendations for chronically infected hosts to take extra precautions to prevent severe illness [60].
@@
CI (typical) $1.0$ $0.25$ $0.25$ $D(i)$
CI (chronic) $0.01$ $0.0$ $0.0$ $D(i)$
HQ $2.0$ $0.25$ $0.25$ 7 – 30
SC $1.0$ $0.5$ 00 dynamic
SD $1.0$ $0.25$ $0.1$ dynamic
Table 1: The micro-distancing parameters (interaction strengths) for the considered NPIs. The micro-duration of CI is limited by the disease progression in the affected agent $i$, $D(i)$. Interaction strengths for CI are set to be significantly lower for chronically infected hosts.
We define SD as a broad behavioural-driven NPI that reduces interaction strengths among individuals, due to stay-at-home orders and other measures (including physical distancing, mask wearing, etc.) adopted during (partial) lockdowns. In other words, the SD compliance level can be interpreted as the fraction of the population that follow the restrictions imposed during (partial) lockdowns. We designed an intervention scenario with a dynamically adjustable SD compliance profile representative of NPIs implemented in many countries during the COVID-19 pandemic [26]. Specifically, we set the SD compliance level to gradually decline annually over the 6-year simulation period, from 50% at the start of the pandemic in 2020 to 20% in the endemic stage from 2023 onwards. In addition, the activation of SD was set to be triggered by a sufficiently high prevalence (e.g., exceeding 200 cases), with SD deactivated once the disease prevalence falls below a certain threshold (e.g., 100 cases). Table 2 and Fig. 5 describe a detailed setup and parametrisation.
@@
CI (typical) $0.7$ $0.7$ $0.7$ $0.7$
CI (chronic) $1.0$ $1.0$ $1.0$ $1.0$
HQ $0.5$ $0.5$ $0.5$ $0.5$
SC (students/teachers) $1.0$ $0.0$ $0.0$ $0.0$
SC (parents) $0.25$ $0.0$ $0.0$ $0.0$
SD $0.5$ $0.4$ $0.3$ $0.2$
Table 2: The macro-distancing parametrisation (population fractions) for the considered NPIs over the 6-year simulation period for SARS-CoV-2. The CI-compliant population fraction is lower for typically infected hosts than for chronically infected hosts. Students/teachers are assumed to fully comply with SC, while parents of school-aged children have a reduced compliance level.
@@ D.1 Vaccination
We simulated a mass vaccination roll-out with a variable daily vaccination rate updated annually, starting from zero (i.e., no available vaccine) in 2020, rising to 0.147% of the entire population per day in 2021 (when vaccines first became available), and then linearly declining to 0.047% of the population per day in 2026 (Fig. 17).
@@ D.2.1 Compound immunity against symptomatic infection
  • $m^{\textrm{c}}(r)$ denotes the peak immunity against symptomatic infection developed from either vaccination or past infection recorded in $r$ against a target variant with genome $s_{T}$ (see Table 3,
@@
Equation 12 shows that $m_{i}^{\textrm{c}}$ reduces linearly over time, governed by two constant rates: immunity waning rate $\epsilon^{c}$ over time, and immunity reduction factor $\tau$ over genetic distance. This assumption matches observations reported in various studies suggesting a linear reduction between (i) COVID-19 vaccine effectiveness and time [69, 70, 71], and (ii) genetic distance and immune evasion [62]. Figure 18 shows a linear regression fitting of vaccine effectiveness over time for different types of vaccines, constructed using supplementary data from [71]. We note that the waning immunity term can be easily adjusted to follow a non-linear reduction, such as exponential decay [72, 73] or based on a gamma distribution [74].
Figure 18: Vaccine effectiveness against infection reduction over time for two widely used mRNA vaccines: (A) Pfizer (BNT162b2), and (B) Moderna (mRNA-1273), plotted and fitted using data from [71]. The equation shown in each plot represents the fitted linear regression for each vaccine, estimating a linear reduction in effectiveness from 2.1% to 2.4% per month.
We then compute the compound immunity for agent $i$ induced by multiple vaccination and/or infection events. The compound immunity against symptomatic infection, $M_{i}^{\textrm{c}}(n,H_{i},s_{x})$, accumulates non-linearly with an upper bound of $1$ (i.e., perfect immunity):
@@ D.2.2 Compound immunity against forward transmission
Following [75, 22, 76], we also consider the immunity component against forward transmission. At time cycle $n$, for an infected agent $i$, the immunity against forward infection from one vaccination/infection event, represented by record $r$, is calculated as follows:
$m_{i}^{\iota}(n,r)=m^{\iota}(r)\Bigl{[}1-\text{min}\Bigl{(}1,\epsilon^{\iota}(n-n_{r})\Bigr{)}\Bigr{]}$ (16)
where $m^{\iota}(r)$ denotes the peak immunity against forward transmission developed from either vaccination or past infection recorded in $r$; rate $\epsilon^{\iota}$ is the forward-transmission immunity waning rate per simulation cycle; and $n_{r}$ is the record cycle of $r$.
We then consider the compound immunity against forward transmission, accounting for all records in $H_{i}$, which follows a non-linear combination of all individual components $m_{i}^{\iota}(n,r)$. Similar to the calculation of the compound immunity against symptomatic infection, we calculate the compound immunity against forward transmission where the upper bound equals $1$ (i.e., perfect immunity):
$M_{i}^{\iota}(n,H_{i})=\text{min}\left(\sqrt{\sum_{r\in H_{i}}\left[m_{i}^{\iota}(n,r)\right]^{2}},1\right)$ (17)

### v3

@@ A.5 Intra-host (within-host) selective pressure
A virus that has less opportunity to develop within-host diversity will be less effective at exploring the fitness landscape. We emulate this by biasing the fitness of mutant sequences generated from a parent sequence. At each time step where we need to generate mutations in the viral sequence associated with an infected host, we first independently generate $M$ candidate sequences and rank them by their fitness. We then sample one sequence from the top $X\leq M$ candidates to be the new sequence. The difference $M-X+1$ specifies the selective pressure strength. For example, if $M=X$, this is equivalent to generating only a single candidate sequence, indicating a weak selective pressure. On the other hand, if $M=100$ and $X=1$, then the selected sequence will be the fittest of 100 randomly generated candidate sequences, indicating a strong selective pressure and bias towards higher fitness. In other words, small values of $X$ imply more efficient exploration of the fitness landscape as might be expected during a chronic infection. This mechanism emulates within-host dynamics that are not explicitly modelled, since each infected agent is associated with only a single viral genome at each simulation cycle. We discussed this distinction in Section 3.
@@
    Randomly generate $M-X+1$ mutated genome sequences by independently selecting and changing $L$ positions in sequence $s$ for each, by substituting the corresponding nucleotide number by a uniformly sampled number between 0 and 333. These generated mutated genome sequences become candidates for the upcoming selection process.
@@ Appendix B Artificial agent-based population
Prior to the simulation, we stochastically generated three distinct artificial populations of anonymous agents (corresponding to layer (B) Demographic shown in Fig 12), used to examine evolutionary and epidemiological dynamics:
@@
We selected these population sizes for two reasons: (i) to investigate the relationship between the population size and phylodynamics, and (ii) to ensure computational feasibility. These considerations are based on the observation that the SARS-CoV-2 variants of concern may have emerged in countries of different population sizes (e.g., South Africa, with a population of approximately 59 million; UK, with a population of around 56 million; and Botswana, with a population over 2.4 million). In addition, our preliminary test simulations revealed that the computational cost increases significantly with the population size (see Section H). In this study, we focus on revealing the representative evolutionary characteristics of SARS-CoV-2, while maintaining computational efficiency by using the three populations listed above. However, given sufficient computational resources, much larger populations (e.g., over one billion population representing countries such as India) may also be generated in order to quantify the impact of population size on phylodynamic and disease dynamics.
The constructed artificial population captures essential demographic characteristics and commuting patterns represented in the latest Australian census and other datasets, as well as agent-to-agent interactions in various social mixing contexts, thus representing the population heterogeneity. Specifically, we generated the artificial population using demographic and travel data sourced from the Australian Bureau of Statistics (ABS) 2021 Census [41], international air traffic reports (detailing incoming passenger flows at Australian airports) [57], and educational registration records (including data on schools and students) [58]. Each agent in the artificial population was assigned multiple demographic attributes, such as age, gender, and residency location, alongside social mixing contexts across various settings: residential (e.g., household, household cluster, neighbourhood and statistical area which maps to a local government area), educational (e.g., classroom/school for agents aged 18 years or younger), and workplace (for agents aged over 18 years). Fig 5 provides a visual representation of the considered social mixing contexts. For each agent, the residential contexts are determined based on residential demographics, while the workplace and educational contexts are assigned based on commuting patterns (commute to work and class/school, depending on the agent’s age group).
@@ C.1 Multi-strain transmission model
The multi-strain transmission model of PhASE TraCE stochastically simulates pathogen transmission between hosts (Epidemic Layer shown in Fig 12). This model introduced several new features to our previous ABM, which was implemented in AMTraC-19 [6, 24, 27, 25]. In this section, we describe how transmission is affected by the phylogenetic model described above.
@@ C.1.1 Susceptible-infectious transition
where $K(s_{j})$ represents the transmissibility of pathogen variant $s_{j}$, proportional to the corresponding basic of effective reproductive number, i.e., $K(s_{j})$ is the fitness of genome $s_{j}$ carried by agent $j$, as defined by Eq 7; $n_{j}$ is the time cycle when agent $j$ started experiencing the onset of infection; and the function $f_{j}(n-n_{j})$ determines agent $j$’s infectivity over time according to the natural history model. At the time cycle $n$, agent $j$’s infectivity is $0<f_{j}(n-n_{j})\leq 1$ with the infectivity peak equal to 1. For uninfected agents, $f(\cdot)=0$. For simplicity, we assume that all variants share the same progression of disease, but differ in their transmissibility (determined by fitness $K$), and that only one variant can be transmitted during the infection process. We note that Eq 8 and Eq 3 in the main manuscript are equivalent.
@@
where $G_{i}(n)$ is the set of social mixing contexts where interactions between agent $i$ and other agents may occur (depending on factors such as weekday or weekend, daytime or nighttime), and $A_{g}\backslash\{i\}$ is the list of all agents in the context $g\in G_{i}(n)$ except agent $i$. We shall expand Eq 9 in later sections to incorporate the effects of non-pharmaceutical interventions (Eq 11) and effects of immunity (Eq 20). We note that Eq 9 and Eq 2 in the main manuscript are equivalent.
Eq 9 defines the infection probability for agent $i$ interacting with other infected agents across all of its social contexts. However, this probability is an aggregate of the interactions and does not identify the specific source of infection, which must be known in order to copy its viral sequence to the newly infected host. We assign the source of infection by constructing a discrete distribution to randomly sample one agent from all potential infected agents $j$ across all social contexts $g\in G_{i}(n)$ to which agent $i$ belongs. The probability of selecting agent $j$ is given by
@@
[…]
@@ C.1.2 Infectious-recovered and recovered-susceptible transitions
Following the natural history model, an infectious agent recovers after a certain recovery period dependent on the host category (see Section G). We assume that both typical and chronically infected hosts develop infectivity fairly quickly from the onset of infection (following a lognormal distribution). However, chronically infected hosts experience a significantly longer recovery than the typical infected hosts. The natural history model describing the disease progression for both host categories is illustrated in Fig  16.
@@ C.2 Non-Pharmaceutical Interventions
The assignment of NPI-compliant agents is determined based on a Bernoulli process with the probability specified by the macro-distancing parameters. While an agent may comply with multiple NPIs, their interaction strengths, $F_{g}(j)$, can only be adjusted to one NPI. We therefore use the parameters for only the first NPI to which an agent complies in the following ordered list: CI, HQ, SD, and SC. The micro and macro-distancing parameters of CI vary depending on the host type, where chronically infected hosts are fully compliant (i.e., macro-distancing level is set to 1.01.01.0), with significantly reduced interaction strengths across all social contexts (Tables 1 and 2). This parametrisation agrees with the public health recommendations for chronically infected hosts to take extra precautions to prevent severe illness [62].
@@
CI (typical) 1.01.01.0 0.250.250.25 0.250.250.25 $D(i)$
CI (chronic) 0.010.010.01 0.00.00.0 0.00.00.0 $D(i)$
HQ 2.02.02.0 0.250.250.25 0.250.250.25 7 – 30
SC 1.01.01.0 0.50.50.5 0 dynamic
SD 1.01.01.0 0.250.250.25 0.10.10.1 dynamic
Table 1: The micro-distancing parameters (interaction strengths) for the considered NPIs. The micro-duration of CI is limited by the disease progression in the affected agent $i$, $D(i)$. Interaction strengths for CI are set to be significantly lower for chronically infected hosts. CI: Case Isolation; HQ: Home Quarantine; SC: School Closure; and SD: Social Distancing.
We define SD as a broad behavioural-driven NPI that reduces interaction strengths among individuals, due to stay-at-home orders and other measures (including physical distancing, mask wearing, etc.) adopted during (partial) lockdowns. In other words, the SD compliance level can be interpreted as the fraction of the population that follow the restrictions imposed during (partial) lockdowns. We designed an intervention scenario with a dynamically adjustable SD compliance profile representative of NPIs implemented in many countries during the COVID-19 pandemic [27]. Specifically, we set the SD compliance level to gradually decline annually over the 6-year simulation period, from 50% at the start of the pandemic in 2020 to 20% in the endemic stage from 2023 onwards. In addition, the activation of SD was set to be triggered by a sufficiently high prevalence (e.g., exceeding 200 cases), with SD deactivated once the disease prevalence falls below a certain threshold (e.g., 100 cases). In our simulations, the deactivation of SD has not been triggered. Fig 5 describes a detailed setup and parameterisation.
@@
CI (typical) 0.70.70.7 0.70.70.7 0.70.70.7 0.70.70.7
CI (chronic) 1.01.01.0 1.01.01.0 1.01.01.0 1.01.01.0
HQ 0.50.50.5 0.50.50.5 0.50.50.5 0.50.50.5
SC (students/teachers) 1.01.01.0 0.00.00.0 0.00.00.0 0.00.00.0
SC (parents) 0.250.250.25 0.00.00.0 0.00.00.0 0.00.00.0
SD 0.50.50.5 0.40.40.4 0.30.30.3 0.20.20.2
Table 2: The macro-distancing parametrisation (population fractions) for the considered NPIs over the 6-year simulation period for SARS-CoV-2. The CI-compliant population fraction is lower for typically infected hosts than for chronically infected hosts. Students/teachers are assumed to fully comply with SC, while parents of school-aged children have a reduced compliance level. CI: Case Isolation; HQ: Home Quarantine; SC: School Closure; and SD: Social Distancing
@@ D.1 Vaccination
We simulated a mass vaccination roll-out with a variable daily vaccination rate updated annually, starting from zero (i.e., no available vaccine) in 2020, rising to 0.147% of the entire population per day in 2021 (when vaccines first became available), and then linearly declining to 0.047% of the population per day in 2026 (Fig 17).
@@ D.2.1 Compound immunity against symptomatic infection
  • $m^{\textrm{c}}(r)$ denotes the peak immunity against symptomatic infection developed from either vaccination or past infection recorded in $r$ against a target variant with genome $s_{T}$ (see Table 3),
@@
Eq 12 shows that $m_{i}^{\textrm{c}}$ reduces linearly over time, governed by two constant rates: immunity waning rate $\epsilon^{c}$ over time, and immunity reduction factor $\tau$ over genetic distance. This assumption matches observations reported in various studies suggesting a linear reduction between (i) COVID-19 vaccine effectiveness and time [71, 72, 73], and (ii) genetic distance and immune evasion [64]. Fig 18 shows a linear regression fitting of vaccine effectiveness over time for different types of vaccines, constructed using supplementary data from [73]. We note that the waning immunity term can be easily adjusted to follow a non-linear reduction, such as exponential decay [74, 75] or based on a gamma distribution [76].
Figure 18: Vaccine effectiveness against infection reduction over time. Two widely used mRNA vaccines are studied: (A) Pfizer (BNT162b2), and (B) Moderna (mRNA-1273), plotted and fitted using data from [73]. The equation shown in each plot represents the fitted linear regression for each vaccine, estimating a linear reduction in effectiveness from 2.1% to 2.4% per month.
We then compute the compound immunity for agent $i$ induced by multiple vaccination and/or infection events. The compound immunity against symptomatic infection, $M_{i}^{\textrm{c}}(n,H_{i},s_{x})$, accumulates non-linearly with an upper bound of 111 (i.e., perfect immunity):
@@
We note that this equation is equivalent to Eq 4 in the main manuscript.
@@
We note that Eq 15 reduces to Eq 14 when $M^{\zeta}_{i}(n,H_{i},s_{x})=M^{\theta}_{i}(n,H_{i},s_{x})$.
@@ D.2.2 Compound immunity against forward transmission
Following [77, 23, 78], we also consider the immunity component against forward transmission. At time cycle $n$, for an infected agent $j$, the immunity against forward infection from one vaccination/infection event, represented by record $r$, is calculated as follows:
$m_{j}^{\iota}(n,r)=m^{\iota}(r)\Bigl{[}1-\text{min}\Bigl{(}1,\epsilon^{\iota}(n-n_{r})\Bigr{)}\Bigr{]}$ (16)
where $m^{\iota}(r)$ denotes the peak immunity against forward transmission developed from either vaccination or past infection recorded in $r$; the rate $\epsilon^{\iota}$ is the forward-transmission immunity waning rate per simulation cycle; and $n_{r}$ is the record cycle of $r$. We assume that $m_{j}^{\iota}(n,r)$ is dependent only on the immunological history of the infection source agent $j$, and is independent of the immunological history of the potentially infected agent.
We then consider the compound immunity against forward transmission, accounting for all records in $H_{i}$, which follows a non-linear combination of all individual components $m_{j}^{\iota}(n,r)$. Similar to the calculation of the compound immunity against symptomatic infection, we calculate the compound immunity against forward transmission where the upper bound equals 111 (i.e., perfect immunity):
$M_{j}^{\iota}(n,H_{j})=\text{min}\left(\sqrt{\sum_{r\in H_{j}}\left[m_{j}^{\iota}(n,r)\right]^{2}},1\right)$ (17)